OpenStudy (anonymous):

Find the solution to 3y^2y'+16x=2xy^3, which is limited when x->+infinite

OpenStudy (anonymous):

i came to y^3=C*e^(x^2)-8

OpenStudy (anonymous):

apply variable separable method, that should get us to the answer

OpenStudy (anonymous):

\[y^3= 8+k*e^{3*(x^2)}\]

OpenStudy (anonymous):

and what does the part about limits, mean? as in, y tends to some value, right? what value?

OpenStudy (anonymous):

it only says that i have to find solution of this equation, which is limited when x goes to \[\infty\]

OpenStudy (anonymous):

and the result should be y=2

OpenStudy (anonymous):

well then, if y is supposed to be limited, the only way that might happen is if the term containing x vanishes as x -> +infinity, right? the only way that might happen is if k=0 and then we might get y=2, if that is the answer.

OpenStudy (anonymous):

oh, y=2 is the answer :P